OpenStudy (anonymous):

Evaluate the limits the lim of abs(2x-4)/2-x as x goes to 2+ , 2- and 2

OpenStudy (helder_edwin):

so we have \[ \Large \lim_{x\to2+}\frac{|2x-4|}{2-x} \] right?

OpenStudy (anonymous):

right

OpenStudy (helder_edwin):

OK. Since \[ x\to2+ \] then x>2 then \[ \Large x>2 \] \[ \Large 2x>4 \] \[ \Large 2x-4>4-4=0 \] then \[ \Large |2x-4|=2x-4 \] do you get this?

OpenStudy (anonymous):

yes

OpenStudy (helder_edwin):

OK, then \[ \Large \lim_{x\to2+}\frac{|2x-4|}{2-x}= \lim_{x\to2+}\frac{2x-4}{2-x}=\lim_{x\to2+}\frac{2(x-2)}{2-x}=-2 \] right?

OpenStudy (anonymous):

right

OpenStudy (helder_edwin):

you can the same with x to 2- using \[ \Large x\to2-\Rightarrow x<2 \] tell me what do you get

OpenStudy (anonymous):

ok the 1 you did at the top what did it come out to be. would the second 1 be -(2x-4) after you take the absolute value away.

OpenStudy (helder_edwin):

Yes! \[ \Large x<2 \] \[ \Large 2x<4 \] \[ \Large 2x-4<0\Rightarrow |2x-4|=-(2x-4) \]

OpenStudy (helder_edwin):

what is the value of the limit on 2-

OpenStudy (anonymous):

is it -2

OpenStudy (helder_edwin):

are you sure?

OpenStudy (anonymous):

I think so but not really sure

OpenStudy (helder_edwin):

let's see \[ \Large \lim_{x\to2-}\frac{|2x-4|}{2-x}=\lim_{x\to2-}\frac{-(2x-4)}{2-x}= \lim_{x\to2-}\frac{-2(x-2)}{2-x} \] \[ \Large =\lim_{x\to2-}\frac{-2}{-1}=2 \]

OpenStudy (anonymous):

O ok so you canceled out the x-2 and the 2-x on top and bottom to get -2 over -1 Right ?

OpenStudy (helder_edwin):

Yes!

OpenStudy (helder_edwin):

now with these two calculations, your final queation: does the limit \[ \Large \lim_{x\to2}\frac{|2x-4|}{2-x} \] exist?

OpenStudy (anonymous):

no Right ?

OpenStudy (helder_edwin):

right!

OpenStudy (anonymous):

thanks you are great

OpenStudy (helder_edwin):

you are welcome